Okay so he question says to; "Rewrite the equation in slope-intercept form. Then identify the slope and y-intercept of the line."
The problem goes as
-2x - y = 3
I was able to rearrange it to go to
-y = 3 + 2x
But, now I have no idea what to do :(

Answer by Alan3354(69443)

Okay so he question says to; "Rewrite the equation in slope-intercept form. Then identify the slope and y-intercept of the line."
The problem goes as
-2x - y = 3
I was able to rearrange it to go to
-y = 3 + 2x
But, now I have no idea what to do :(
-----------------
You were almost done.
-y = 3 + 2x
y = -2x - 3
Slope m = -2
y-intercept = -3